OpenStudy (anonymous):

Solve. y/5 = 5 5 * 5 = y

OpenStudy (anonymous):

right

OpenStudy (ivettef365):

right so y = ??

OpenStudy (anonymous):

y = ????????????????????????????????????????????????????????????????25?????????????????????????????

OpenStudy (texaschic101):

to get rid of a fraction, you multiply everything by the denominator. y/5 = 5 (multiply everything by 5) (5) y/5 = (5)5 y = 5 * 5 y = 25

OpenStudy (anonymous):

gj

OpenStudy (texaschic101):

thanks :)